OpenStudy (anonymous):

Simplify: (4x-y)(4x+y)

OpenStudy (anonymous):

You need to foil. F = 4x * 4x = 16x^2 O = 4x* y = 4xy I = -y * 4x = -4xy L = -y * y = -y^2 Add like terms and you get 16x^2 - y^2 Hint... if you have the exact same first terms and last terms... (4x y)(4x y) and you have one + sign and one - sign you can get the answer by multiplying the F and the L only because the O and I will ALWAYS cancel out. Be careful though, you can only do that when they are identical with one + sign and one - sign.

OpenStudy (anonymous):

ok thanks